13. Tartaric acid, H:C:H:04, is a diprotic acid with Kal = 9.20*10* and K2= 4.31x10. Determine the pH and the concentrations please be very clear and detailed on how you got the answer and why for a thumbs up
Answer #1

Dipostic acid (H2 C4 HAOA) : 0.261 m и, са и, од исцидо, Сва) + н* ta) 0-260 Ба, д.2x104 н от - Х + x О. 26) - ка. - См] [иat equilibrium, [12 CQ HA Oq] = 0.261-1 = 0.261 - 1.549 x 10-2 = 0.261 -0.01549 = 0.24551 . [4₂ KG Hq Oq J = 0.24551 MpH = -log Cut] = -log (1.549 * 182) = 1.80 IPH = 180 Now other reaction. + 47 ka = 4-31x 10-s HEAHOO I 1.549810-2 c -k E 1.54